# 3DGraphLLM: Combining Semantic Graphs and Large Language Models for 3D Scene Understanding

### v2

@@ 4.1 Experimental Results
Figure 5: Dependence of inference speed and visual grounding accuracy on the number of nearest neighbors in the object subgraph. This experiment utilizes the RioRefer dataset along with GT instance segmentation.
Neural network segmentation imperfections impact both the quality of object embeddings generated by the 3D Object Encoder and the embeddings of semantic relations between objects. We perform a PCA analysis of Uni3D object embeddings and VL-SAT relation embeddings, comparing results for ScanNet training scenes using GT instance segmentation and Mask3D instance segmentation (see Figure 4).
Our analysis shows that, with the standard selection of nearest neighbors, the relation embeddings differ significantly between GT and Mask3D three-dimensional masks. When applying the minimal distance filter, the similarity of VL-SAT relation embeddings significantly increases between GT and Mask3D instance segmentation.
By examining the minimum distance between neighboring objects, we observed that duplicate objects were often selected as neighbors. To address this issue, we introduced a minimum distance filter of 1 cm between neighboring objects, which made the relationship embeddings from GT masks and Mask3D results more consistent. Additionally, applying this filter improved performance on the visual grounding task, as shown in Table 4.
We also experimented with adding an NMS filter to remove duplicates among the vertices that an object may be associated with, with a threshold of $IoU=0.99$. The results in Table 4 show that adding the filter allows for further improvement of the grounding quality.
Table 5: Ablation study on spatial relation module on RioRefer dataset (GT Instance segmentation).
Methods Edge Number Spatial relation Acc@0.5↑
3DGraphLLM 0 $42.6$
3DGraphLLM 2 $48.9(+6.3\%)$
3DGraphLLM 2 $\mathbf{50.1}(+7.5\%)$
Number of nearest neighbors.
We conducted an experiment to examine how the number of nearest neighbors affects the quality of visual grounding and the speed of model inference, as adding more connections increases the number of tokens used to describe each object.
@@
This experiment was performed using ground-truth scene segmentation and the RioRefer dataset (Miyanishi et al., 2024), as this setup provides the highest quality embeddings for semantic relations between objects. We vary the number of nearest neighbors in powers of two, capping it at 5 due to GPU memory constraints during training. As shown in Figure  5, increasing the number of nearest neighbors enhances visual grounding quality with a slight increase in inference time.
Spatial relations. Previous research (Wang et al., 2023b; Huang et al., 2023) has shown that incorporating spatial relationships between objects, represented by 3D coordinates of their bounding boxes, can improve performance in visual grounding tasks. We attempted to integrate spatial relations into our method by using the output of the spatial transformer as the final token in the relation triplets between an object and its nearest neighbors (i.e., a triplet $(F_{i}^{v},F_{ij_{k}}^{e},F_{j_{k}}^{\text{rel}})$, where $F_{j_{k}}^{\text{rel}}$ represents the output of the Chat3Dv2 spatial relation module (Huang et al., 2023)). However, as shown in Table 5, our experiments did not find this approach effective for learning a graph representation of a scene.
@@ 5 Conclusion
In this paper, we propose a new learnable approach to using a 3D semantic scene graph for a large language model solving the 3D vision-language tasks. Detailed experiments demonstrate the effectiveness of this approach, which explicitly takes into account semantic relations between objects represented as 3D point clouds. Our approach, called 3DGraphLLM, demonstrated state-of-the-art quality on popular ScanRefer, Multi3DRefer, and Scan2Cap datasets.
@@
For further development of the work, it seems appropriate to search for the methods to reduce token usage for encoding object relationships in our graph representation. Another important aspect for further work is the creation of methods for generating semantic relations between objects that are robust to imperfections in the instance segmentation of the scene point cloud.
Appendix A Common Failure Cases
We illustrate the most common failure cases of 3DGraphLLM related to spatial relationships in Figure 6.
Figure 6: Common failure cases of 3DGraphLLM related to spatial relationships. Left: In the ScanQA dataset, 3DGraphLLM incorrectly identifies the front/back and left.right directions relative to the observer. Right: In the ScanRefer dataset, 3DGraphLLM confuses left and right. The GT object is highlighted in green, and the 3DGraphLLM prediction is highlighted in red.
Appendix B Functional Queries
We illustrate the ability of 3DGraphLLM to leverage common sense knowledge in its responses to question types not present in the training dataset in Figure 7.
@@ Appendix B Functional Queries
Figure 7: Functional queries about the room and objects to the 3DGraphLLM. Left: 3DGraphLLM is capable of answering questions about functional properties of the room and its room type. Right: 3DGraphLLM is capable of answering questions about the functional properties of objects in a room.

### v3

@@ 4.2 Ablation Studies
Figure 4: Dependence of inference speed and visual grounding quality on the number of nearest neighbors in the object subgraph. This experiment utilizes the GT instance segmentation.
The analysis of objects selected as nearest neighbors reveals a high number of duplicate objects among the chosen neighbors. To address this issue, we propose two filters. First, we add an NMS filter to remove duplicates between the potential neighbors for an object, using a threshold of $IoU=0.99$. Second, we introduce a minimum distance filter of 1 cm to the nearest neighbor to prevent selecting duplicates of the original object as its neighbors.
Adding the NMS filter improves the performance of the visual grounding task when using Mask3D instance segmentation (see  Tab. 5). The additional minimum distance filter further enhances visual grounding quality. The combination of filters is also effective for OneFormer3D [33] scene instance segmentation, as shown in  Tab. 5.
@@
Number of nearest neighbors. We examine how the number of nearest neighbors affects the quality of visual grounding and the speed of model inference, as adding more connections increases the number of tokens used to describe each object. This experiment was performed using ground-truth scene segmentation, as this setup provides the highest quality embeddings for semantic relations between objects. We vary the number of nearest neighbors in powers of two, capping it at 4 due to GPU memory constraints during training. As shown in  Fig. 4, increasing the number of nearest neighbors enhances visual grounding quality with a slight increase in inference time.
Subgraph representation. In our work, we use an object-centric graph representation, where relationships between objects are represented as triplets $\{F_{N}^{v},F_{Nk_{1}}^{e},F_{k_{1}}^{v}\}$. We conduct an experiment in which we remove duplicate vertex tokens from the subgraph-based object description. As a result, object $N$ is described by the following sequence: $\{<OBJN>F_{N}^{2d},F_{N}^{v},F_{Nk_{1}}^{e},F_{Nk_{2}}^{e}\}$. We do not perform the pretraining phase on GT instance segmentation in this experiment.  Tab. 8 shows that the object-centric graph representation using triplets improves the performance of the visual grounding task.
We include additional experimental results from ablation studies on scene captioning and visual question answering tasks in the Supplementary Materials.
@@ 5 Conclusion
In this paper, we propose a new learnable approach to using a 3D semantic scene graph for a large language model to solve 3D vision-language tasks. Detailed experiments demonstrate the effectiveness of this approach, which explicitly takes into account semantic relations between objects represented as 3D point clouds. Our method, called 3DGraphLLM, surpasses the baseline approach without semantic relationships on popular ScanRefer, Multi3DRefer, Scan2Cap, ScanQA, and SQA3D datasets. Moreover, 3DGraphLLM achieves state-of-the-art performance in the object grounding task, matching the quality of methods that require five times more inference time.
@@
For further development of the work, it seems appropriate to search for methods to reduce token usage for encoding object relationships in our graph representation. Another important aspect for further work is the creation of methods for generating semantic relations between objects that are robust to imperfections in the instance segmentation of the scene point cloud.
Acknowledgments
The study was supported by the Ministry of Economic Development of the Russian Federation (agreement with MIPT No. 139-15-2025-013, dated June 20, 2025, IGK  [^00] C313925P4B0002).
Appendix A Ablation Study. Number of Nearest Neighbors
We investigate how increasing the number of nearest neighbors affects the quality of scene description and question answering tasks. The number of nearest neighbors varies from 0 to 4.
When no nearest neighbors are used, the model operates as a baseline Chat-Scene approach, representing the scene as a list of objects without semantic relationships. The maximum number of nearest neighbors that can be added for each object is four, constrained by GPU memory limitations during model training. In these experiments, we use GT instance segmentation to eliminate errors in graph construction that could otherwise impact performance on 3D vision-language tasks. The base LLM used in this study is LLAMA3-8B-Instruct.
As shown in  Fig. 5, increasing the number of nearest neighbors improves the quality of object descriptions in the scene, while causing only a slight increase in generation time. Notably, in dense scene captioning tasks, inference speed depends not only on the number of tokens used to represent the scene but also on the length of the generated description. Therefore, we observe only a small increase in generation time when comparing graphs with two and four nearest neighbors.
Fig. 6 and Fig. 7 show that semantic relationships between objects enhance performance in question answering tasks on the ScanQA and SQA3D datasets. However, for question answering tasks, the optimal number of nearest neighbors is 2, as increasing it to 4 leads to a drop in performance. The impact of semantic edges is harder to assess in question answering tasks than in visual grounding or object description tasks, since some question types do not require knowledge of object spatial relationships.
For further experiments, we select 2 nearest neighbors, as it provides the best trade-off between performance gains and computational complexity across all three tasks.
Figure 5: Dependence of inference speed and dense scene captioning quality on the number of nearest neighbors in the object subgraph. This experiment utilizes the GT instance segmentation.
Figure 6: Dependence of inference speed and question answering quality on ScanQA dataset on the number of nearest neighbors in the object subgraph. This experiment utilizes the GT instance segmentation.
Instance Number Minimal ScanRefer Multi3DRefer Scan2Cap ScanQA Sqa3D
Methods segmentation of edges distance, cm Acc@0.5↑ F1@0.5↑ C@0.5↑ B-4@0.5↑ C↑ B-4↑ EM↑
3DGraphLLM-0 GT 0 - 61.5 64.4 89.5 40.1 83.1 15.6 55.2
3DGraphLLM-2 GT 2 0 66.9 69.9 95.2 42.7 88.6 16.2 56.3
3DGraphLLM-0 Mask3D 0 - 52.0 55.1 80.0 37.5 84.0 15.8 53.8
3DGraphLLM-2 Mask3D 2 0 55.6 58.2 80.8 36.4 85.7 15.1 56.0
3DGraphLLM-2 Mask3D (+ NMS) 2 0 55.7 58.6 82.3 36.8 86.2 16.0 56.2
3DGraphLLM-2 Mask3D (+ NMS) 2 1 56.2 58.7 82.9 37.3 85.4 15.1 55.6
3DGraphLLM-0 OneFormer3D 0 - 50.0 52.8 73.5 34.3 87.3 16.5 53.8
3DGraphLLM-2 OneFormer3D 2 0 52.8 55.8 70.2 32.7 83.3 15.0 55.0
3DGraphLLM-2 OneFormer3D (+NMS) 2 1 54.6 57.2 72.4 33.0 81.3 12.9 55.0
Table 7: Ablation study on semantic edge role depending on quality of instance segmentation.
Figure 7: Dependence of inference speed and question answering quality on SQA3D dataset on the number of nearest neighbors in the object subgraph. This experiment utilizes the GT instance segmentation.
Appendix B Ablation Study. Quality of Instance Segmentation.
For the graph, where the vertices contain objects obtained from 3D instance segmentation, we observe a consistent improvement in the performance of all three 3D Vision-Language tasks. When transitioning from GT segmentation to a noisy graph composed of vertices obtained using Mask3D, we also observe improvements in metrics for all three tasks (see  Tab. 7). However, this improvement is less pronounced compared to GT instance segmentation.
We compare the graphs of nearest neighbors obtained from GT instance segmentation and Mask3D instance segmentation. The analysis shows that Mask3D instance segmentation contains a large number of object duplicates, as the scene is always divided into N=100 segments. The presence of object duplicates among the neighbors leads to a reduction in useful information about the object’s environment in its subgraph. To address the duplicates in the vertices of the subgraphs, we use two filters. The NMS filter with an IoU threshold of 0.99 removes duplicate objects from the neighbors. The minimum distance filter between the centers of the object point clouds excludes the object’s own duplicates from its neighbors.
Tab. 7 shows that adding these filters consistently improves the performance of visual grounding and object description tasks for the graph obtained through Mask3D instance segmentation. Since we expect an effect from adding semantic edges specifically for these tasks, we keep this filter in further experiments.
We also experiment with different methods for instance segmentation to create scene graph vertices. We use another method for instance segmentation, OneFormer3D, filtering out vertices with confidence <0.1<0.1. We observe that for scene graphs with such vertices, semantic edges improve the performance of the visual grounding task. At the same time, the combination of nearest-neighbor filters proves effective for this type of scene segmentation, increasing the performance of object grounding and scene captioning tasks. However, since OneFormer3D showed worse results with these hyperparameters compared to Mask3D, and other baseline methods use Mask3D for object proposals, we chose Mask3D for the final version of the pipeline.
Appendix C Ablation Study. Subgraph Representation.
We explore the possibility of further flattening the graph by replacing relationship triplets with a sequence of semantic edges. As a result, object $N$ is described by the following sequence: $\{<OBJN>F_{N}^{2d},F_{N}^{v},F_{Nk_{1}}^{e},F_{Nk_{2}}^{e}\}$.
Instance Relations Number ScanRefer Multi3DRef Scan2Cap ScanQA Sqa3D
Methods segmentation as triplets of edges Acc@0.5↑ F1@0.5↑ C@0.5↑ B-4@0.5↑ C↑ B-4↑ EM↑
3DGraphLLM-0 Mask3D 0 52.0 55.1 80.0 37.5 84.0 15.8 53.8
3DGraphLLM-2 Mask3D 2 54.2 56.3 87.2 39.3 85.6 15.1 54.6
3DGraphLLM-2 Mask3D 2 54.3 57.3 85.6 39.6 87.4 14.9 54.5
Table 8: Ablation study on subgraph representation.
@@ Acknowledgments
Figure 8: Functional queries about the room and objects to the 3DGraphLLM. 3DGraphLLM is capable of answering questions about the functional properties of the room and its room type as well as discerning the functional properties of objects in a room.
Figure 9: Common failure cases of 3DGraphLLM related to spatial relationships. In the question answering task, 3DGraphLLM incorrectly identifies the front/back and left/right directions relative to the observer. In the visual grounding task, 3DGraphLLM confuses left and right. The GT object is highlighted in green, and the 3DGraphLLM prediction is highlighted in red. In the object captioning task, 3DGraphLLM uses a spatial reference not mentioned in the GT descriptions. However, the description is correct qualitatively.
Tab. 8 shows that representing relationships as triplets improves the performance of the object grounding task. The object-centered representation of relationships improves question answering performance on ScanQA according to the CIDEr metric, while the BLEU-4 metric remains on par with the flat graph representation as a sequence of relationships. For the SQA3D dataset, both approaches yield comparable results. In the dense scene captioning task, we observe a decrease in the CIDEr metric but an improvement in the Scan2Cap metric. However, as shown in  Appendix E, n-gram-based metrics may produce unreliable results when evaluating text generated by LLMs. Considering this, and the fact that the impact of semantic edges is most interpretable in the object grounding task, we represent relationships as triplets in subsequent experiments.
Appendix D Functional Queries
We illustrate the ability of 3DGraphLLM to leverage common sense knowledge in its responses to question types not present in the training dataset in  Fig. 8.
Appendix E Common Failure Cases
We illustrate the most common failure cases of 3DGraphLLM related to spatial relationships in  Fig. 9.
It is important to note that the quality metrics in the Scan2Cap, ScanQA, and SQA3D benchmarks are based on n-gram-based metrics comparing generated answers with reference ones, such as BLEU-1, BLEU-2, BLEU-3, BLEU-4, CIDEr, ROUGE-L, METEOR. The Exact Matching (EM) metric compares the exact match of the answer with the GT answer. The drawback of these metrics is that if an object description or answer to a question contains spatial relationships not present in the reference descriptions, it leads to a decrease in the score. Additionally, these metrics are unable to adequately evaluate LLM responses, considering the richness of formulations and the freedom to choose visual and spatial properties of an object that may be mentioned by the model. These responses represent a special type of ”failure cases,” illustrated in  Fig. 9 on the right. For this type, the object description or answer to the question is correct from a qualitative point of view but shows zero value according to the metric CIDEr.
Appendix F Scalability with Number of Scene Objects
To evaluate scalability, we analyze how memory usage and inference time vary with the number of objects in a scene. The maximum number of objects considered corresponds to the highest counts observed in the ScanNet and 3RScan datasets. As shown in Tab. 9, both memory usage and inference time increase gradually as the number of objects in the scene grows. This demonstrates that while resource consumption scales with object count, the growth remains manageable and does not compromise the method’s practical applicability within real-world scene configurations.
Number of Objects Memory Usage (Gb) Inference Time for 1 token (s)
10 23 0.08
50 28 0.14
100 35 0.23
Table 9: Performance metrics for varying number of objects in a 3D scene.